OpenStudy (anonymous):

20% of what number is 27

OpenStudy (anonymous):

20% of what number is 27

OpenStudy (whpalmer4):

20% of x = 27 20/100 * x = 27 Multiply both sides by 100 20 * x = 2700 Divide both sides by 20 x = 2700 / 20 = ?

OpenStudy (whpalmer4):

Or, recognize that 5 * 20 % = 100 %. if 27 is 20% of our number, 100% of our number must be 5*27.

OpenStudy (whpalmer4):

Yet another equivalent way: 20% = 20/100 = 0.20. If 0.20*x = 27, divide both sides by 0.20 and get x = 27/0.20 =

OpenStudy (whpalmer4):

Pick the method that seems easiest for you to understand, and practice with it until it becomes second nature.